40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Dec 7 11:25:31 2020, gautam, Update, SUS, MC1 suspension glitchy again 
    Reply  Wed Dec 9 16:22:57 2020, gautam, Update, SUS, Yet another round of Sat. Box. switcharoo WFS2.pngWFS_lineNoise.pngWFSchar.pdf
       Reply  Sun Jan 3 16:26:06 2021, Koji, Update, SUS, IMC WFS check (Yet another round of Sat. Box. switcharoo) Screen_Shot_2021-01-03_at_17.14.57.png
    Reply  Thu May 13 11:55:04 2021, Anchal, Paco, Update, SUS, MC1 suspension misbehaving Screenshot_from_2021-05-13_09-50-24.pngMC1_Glitches_Invest2.pdf
       Reply  Thu May 13 19:38:54 2021, Anchal, Update, SUS, MC1 Satellite Amplifier Debugged MC1_UL_Channel_Fixed.png
          Reply  Mon May 24 19:14:15 2021, Anchal, Paco, Summary, SUS, MC1 Free Swing Test set to trigger 
             Reply  Tue May 25 10:22:16 2021, Anchal, Paco, Summary, SUS, MC1 new input matrix calculated and uploaded SUS_Input_Matrix_Diagonalization.pdf
          Reply  Thu Jun 17 11:45:42 2021, Anchal, Paco, Update, SUS, MC1 Gave trouble again SummaryScreenShot.pngMC1_LL_SENSOR_DEAD.png
             Reply  Thu Jun 17 16:37:23 2021, Anchal, Paco, Update, SUS, c1susaux computer rebooted 
                Reply  Tue Jun 22 11:56:16 2021, Anchal, Paco, Update, SUS, ADC/Slow channels issues CDS_FE_Status.png
                   Reply  Tue Jun 22 16:52:28 2021, Paco, Update, SUS, ADC/Slow channels issues shake_and_damp.png
                      Reply  Wed Jun 23 09:05:02 2021, Anchal, Update, SUS, MC lock acquired back again 
                         Reply  Thu Jun 24 16:40:37 2021, Koji, Update, SUS, MC lock acquired back again P_20210624_163641_1.jpg
Message ID: 16139     Entry time: Thu May 13 19:38:54 2021     In reply to: 16138     Reply to this: 16157   16209
Author: Anchal 
Type: Update 
Category: SUS 
Subject: MC1 Satellite Amplifier Debugged 

[Anchal Koji]

Koji and I did a few tests with an OSEM emulator on the satellite amplifier box used for MC1 which is housed on 1X4. This sat box unit is S2100029 D1002812 that was recently characterized by me 15803. We found that the differential output driver chip AD8672ARZ U2A section for the UL PD was not working properly and had a fluctuating offset at no input current from the PD. This was the cause of the ordeal of the morning. The chip was replaced with a new one from our stock. The preliminary test with the OSEM emulator showed that the channel has the correct DC value.

In further testing of the board, we found that the channel 8 LED driver was not working properly. Although this channel is never used in our current cable convention, it might be used later in the future. In the quest of debugging the issue there, we replaced AD8672ARZ at U1 on channel 8. This did not solve the issue. So we opened the front panel and as we flipped the board, we found that the solder blob shorted the legs of the transistor Q1 2N3904. This was replaced and the test with the LED out and GND shorted indicated that the channel is now properly providing a constant current of 35mA (5V at the monitor out).


After the debugging, the UL channel became the least noisy among the OSEM channels! Mode cleaner was able to lock and maintain it.

We should redo the MC1 input matrix optimization and the coil balancing afterward as we did everything based on the noisy UL OSEM values.

Attachment 1: MC1_UL_Channel_Fixed.png  1.248 MB  Uploaded Thu May 13 21:12:39 2021  | Hide | Hide all
MC1_UL_Channel_Fixed.png
ELOG V3.1.3-